### v1

@@ 2.3 Representation of grain boundaries
GBs are represented as inclusions with different properties compared to the surrounding grains. We use MS calculations to define three GB properties: (1) width (2) second-order elastic constants (SOECs), and (3) lattice energy. The width of the GB was determined using a two-class Gaussian Mixture Model clustering of the atomistic configuration as explained by Mishra et al., [2024]. To determine the SOECs of the GB region, we performed affine deformation of a bicrystal with 3D periodic boundary conditions. A prescribed set of lattice strains () was applied to an energy minimized bi-crystal under the assumption of general elastic anisotropy. Per-atom energies for the atoms in the GB region were used to calculate the energy density in the GB region as a function of the applied strain. The elastic stiffness coefficients were determined from these energy density-strain data. We assume that the strain in the bi-crystal is homogeneous and equal to the applied strain. The Lamé constants are determined from the SOECs using the Voigt average (See Supplementary Figure 1).
We perturb the lattice energy in the GB region using an additional term that accounts for the difference in atomic structure and its effect on the evolution of the dislocation core. This additional energy barrier is relevant to the resulting dislocation-GB interaction; however, it cannot be determined analytically but it can be estimated from atomic scale simulations. Previous studies have estimated the energy barrier using nudged elastic band (NEB) method [Tsuru et al.,, 2009] and MD simulations [Sangid et al.,, 2011, Li et al.,, 2020, 2022]. In Sangid et al., [2011] and Li et al., [2020], the energy barrier and GB energy are correlated using a power law equation; however, the number of data points used are not enough to create a robust model. Thus, in this work, we take a different approach and use the GB energy to scale the lattice energy in the bulk grains (see Equation 22) to define the perturbation on the lattice energy in the GB region as
where is a fitting constant, is the GB energy, is the unstable stacking fault energy, is the shear modulus of the grain and is the shear modulus of the GB. This term acts as a perturbation of the -surface in the bulk grain to account for the disordered nature of the GB with respect to the bulk crystal so the total energy at the GB region is calculated as . To align with previous work, this term could be modified as where and are fitting parameters [Sangid et al.,, 2011, Li et al.,, 2020]. We highlight again that the energy barrier to dislocation glide within a GB due to the atomic disorder is still an open research question, and is expected to have strong dependence on the GB structure. Thus, these equations are representative of this effect, but may need to be redefined to GB structures not considered in this paper. In particular, specialized GB structures such as cube-on-cube orientation or coherent twin boundaries with low values of GB energy , are likely examples of cases in which different descriptions of this energy barrier may be necessary.
2.4 Material properties and simulation setup
@@
In previous work [Dang et al.,, 2025], dislocation-GB interactions were studied with MD for several different symmetric tilt GBs in Cu including both stable and metastable GB configurations. From this database of GB structures [Dang et al.,, 2025], we selected two 110 symmetric tilt GBs and their respective metastable states to consider in this paper. The first GB is a low angle GB (LAGB) with (1,1,6) GB plane and a misorientation angle of 26.5∘ and the second GB is a high angle GB (HAGB) with (5,5,2) GB plane and a misorientation angle of 148.4∘. These GBs have different GB widths, GB energies, elastic constants and GB-dislocation reactions as predicted with MD, see Table 1. The Lamé constants of the Cu grains surrounding the GB region are and and mJ/m2. For each GB, the minimum energy structure plus two higher metastable structures were selected for this study. We note that we were unable to determine the SOECs for some metastable GBs as they were unstable under some deformation modes. Therefore, the Lamé constants and GB width of the minimum energy GB structure is used also for the metastable cases.
@@
A domain of 256x4x128 is used with a grid size that corresponds to the Burgers vector norm of a screw dislocation in Cu ( Å). Periodic boundary conditions are used in the three directions and the domain size is chosen such that the periodic interactions are minimized. An inclusion, representing the GB region, is located in the center of the domain and the position of the initial dislocation dipole is presented in Figure 2. A screw dislocation dipole is placed in the left grain, with the left monopole placed at 63 grid points from the domain boundary with a distance of 50 grid points between the dislocations. The domain is subjected to three different applied shear stresses MPa, resulting in a different reaction for a given GB energy value as shown in Table 1. The dislocation mobility is set to with .
3 Results and discussions
The reactions in Table 1 are reproduced using the PFDD model by setting for each type of GB. The values used in the following simulations for the LAGB and HAGB are 0.04 and 0.086, respectively. As the elastic constants and energies are fixed for a given GB, Equation 24 may be redefined as
@@ 3 Results and discussions
The value of is shown in Table 2 for each GB. Based on the reactions at different applied stress, the six GBs can be divided into three categories as shown in Table 2. One can notice the value of is close within each category. If is between 0.213 and 0.304, the GB-dislocation reactions are (A/P, T, T), if is between 0.361 and 0.41, the GB-dislocation reactions are (A/P, A/P, T), and if is equal or higher than 0.442, the GB-dislocation reactions are (A/P, A/P, A/P). The values roughly correlate with strain-gradient metrics from the strain functional descriptors used in the two-class Gaussian Mixture Model clustering in [Mishra et al.,, 2024]. More details are provided in the supplemental information.
@@
| LAGB | 751 | 0.04 | 0.213 | A/P | T | T |
| LAGB | 1003 | 0.04 | 0.304 | A/P | T (C1) | T |
| LAGB | 1228 | 0.04 | 0.386 | A/P | A/P (C2) | T |
| HAGB | 847 | 0.086 | 0.361 | A/P | A/P (C3) | T (C4) |
| HAGB | 938 | 0.086 | 0.410 | A/P | A/P | T |
| HAGB | 998 | 0.086 | 0.442 | A/P | A/P | A/P |
@@
Among the cases shown in Table 1, four different cases are analyzed here in more detail: (C1) LAGB with GB energy mJ/m2 under an applied stress of MPa, (C2) LAGB with GB energy mJ/m2 under an applied stress of MPa, (C3) HAGB with GB energy mJ/m2 under an applied stress of MPa and (C4) HAGB with GB energy mJ/m2 under an applied stress of MPa. Note that the other cases have been simulated and the GB-dislocation reactions from MD are reproduced, see Supplementary Figures 2-15. C1, C2 and C3 have different energies and same applied stress while C3 and C4 are the same GB with different applied stresses. for C1, C2 and C3/C4 is 0.304, 0.386 and 0.361, respectively (see Table 2). To the authors knowledge, there is no way to compute analytically the values of , for the work presented here the values are chosen to reproduce the MD reactions.
@@
Figures 3, 4, 5 and 6 show the change of the disregistry and stress evolution at different time steps for cases C1, C2, C3 and C4. The disregistry is computed as [Shen and Wang,, 2004]
@@
where is the Burgers vector direction of the initial perfect dislocation. The dislocation transmits in cases C1 and C4, and it is pinned in cases C2 and C3 as the dislocation stays blocked inside (Figure 4) or next (Figure 5) to the GB region. Three main observations need to be highlighted, (i) transmission is easier for LAGB which is in accordance with MD simulations and experimental observations [Nieto-Valeiras et al.,, 2024, Adams et al.,, 2019], (ii) the change in GB structure between C1 and C2 is reflected in the increase of GB energy and this additional GB energy results in an increased resistance to the dislocation and results in the dislocation being pinned, and (iii) the change in the interaction between cases C3 and C4 from being pinned to transmitted is a direct effect of the applied stress. Also, in C4 the dislocation evolves faster due to the higher applied stress of 750 MPa.
@@
The dislocation disregistry change during the interaction is shown in Figure 7. Moreover, the change in the stacking fault width is more clear by plotting its derivative along the slip direction, see Figure 8. These detailed views of the disregistry are shown at the same time steps in Figures 3, 4, 5 and 6. The initial value of the stacking fault width in the four cases is 4 or 5 grid points that correspond to 1nm which is in agreement with previous studies [Jin et al.,, 2006]. For the cases where the dislocation transmits across the GB, there is a difference in the stacking fault width during the interaction. In case C1, the stacking fault width remains constant with values between 4 and 5 grid points, see Figures 3, 7a and 8. While in case C4, the stacking fault width decreases to 2 grid points (Figure 6b) and comes back to its initial shape (Figure 6f). Something similar happens in the cases that get pinned. In case C2, the dislocation gets pinned with the trailing partial being blocked at the interface between the left grain and the GB region and the leading partial stays inside the GB region, also the stacking fault width changes slightly to 3 grid points as shown in Figures 7b and 8. In case C3, as the dislocation approaches the GB and gets pinned the stacking fault width decreases to half of its original width from 4 to 2 grid points, see Figures 7c and 8. Due to this difference, the interaction in case C2 may be categorized as absorption instead of pinning. Note that absorption is a challenging interaction to model as the interaction can have different scenarios, i.e., dislocations can glide and re-emit later, or the dislocation can change the GB structure leading to nucleation of new dislocations [Jin et al.,, 2006, Dang et al.,, 2024, Liang et al.,, 2019, Wang,, 2015]. However, modeling absorption and subsequent dislocation glide along the GB plane is not available in the present PFDD model but could be considered by adding existing glide and cross slip PFDD extensions [Fey et al.,, 2022, Wise et al.,, 2025].
@@
In summary, the variation in the stacking fault width is higher in C3 and C4 as shown in Figures 7 and 8. As is added to the total energy, it acts as an additional attractive force between the two partial dislocations which explains the change in the disregistry. Hence, the difference between the stacking fault width in C1, C2, C3 and C4 is proportional to despite of the higher applied stress in C4. In the Supplementary Material, is plotted against averaged strain gradient metrics showing good correlation with both, net strain gradient and gradient on uniaxial strain metrics computed using MS [Mishra et al.,, 2024]. Similar changes in the stacking fault width have been reported in MD simulations by Jin et al., [2006] (cf. Figure 2) where the partial dislocation is constricted to an almost perfect dislocation at the GB before transmission or absorption of the partial dislocation, and Dang et al., [2024] (cf. Figure 2) where the partial dislocation stacking fault width reduces to almost half of its original width when the dislocation is pinned.
@@ 4 Conclusion
A new PFDD formulation that couples phase field and micromechanical solvers is presented. This new formulation allows for the study of heterogeneous materials with different stiffness tensor by solving the Lippmann-Schwinger equation. The total strain obtained from the Lippmann-Schwinger equation is one of the terms in the elastic energy density used within the time-dependent Ginzburg-Landau (TDGL) kinetic equation.
In this work, GB-dislocation interactions are studied by representing GBs as inclusions with an specific width obtained from MS simulations [Mishra et al.,, 2024]. Two additional parameters are added to describe GBs, (i) an elastic energy which is considered by computing a different GB elastic stiffness tensor and (ii) a lattice energy that represents the effect of the disordered atomic structure within the GB region. Two different groups of GBs are studied including stable and metastable states: LAGB with GB plane (1,1,6) and misorientation angle of 27∘ and HAGB with GB plane (5,5,2) and misorientation angle of 178∘. The GB-dislocations interactions predicted using PFDD show excellent agreement with the MD results. However, these results depend on the calibration parameter which varies as a function of the misorientation. We presented an analysis on the stacking fault width, the initial width and reduction of the width during the interaction which shows good agreement with MD simulations [Jin et al.,, 2006, Dang et al.,, 2024]. Finally, by adding an additional lattice energy for the GB region proportional to , we can account for the dependency of GB-dislocation interaction on the GB energy or GB structure.
In contrast to the sharp interface method [Zeng et al.,, 2016, Ma et al.,, 2022], this new methodology opens the door to model more general GB-dislocation interactions. Future work will be focused on the improvement of the GB description and the inclusion of other interactions such as GB absorption and glide of the dislocation across the GB. Also, other GB-dislocation interactions such as reflection should be considered.

### v2

@@ 2.3 Representation of grain boundaries
In previous work [Dang et al., 2025], dislocation-GB interactions were studied with MD for several different symmetric tilt GBs in Cu including both stable and metastable GB configurations. From this database of GB structures [Dang et al., 2025], we selected two 110 symmetric tilt GBs and their respective metastable states to consider in this paper. The first GB is a low to medium misorientation angle GB (LAGB) with (1,1,6) GB plane and a misorientation angle of 26.5∘ and the second GB is a high misorientation angle GB (HAGB) with (5,5,2) GB plane and a misorientation angle of 148.4∘. In this work, GBs are represented as inclusions with different properties compared to the surrounding grains. We use MS calculations to define three GB properties: (1) width (2) second-order elastic constants (SOECs), and (3) GB energy (GBE). The width of the GB was determined using a two-class Gaussian Mixture Model clustering of the atomistic configuration as explained by Mishra et al. [2024]. Figure 2 shows the symmetric tilt GBE as a function of the misorientation angle, highlighting the GBs selected for this work in boxes. The minimum energy or stable state correspond to the lowest GBE value for a given misorientation angle, while the metastable states have higher GBE values. The GBE of the LAGB and HAGB are colored in blue and orange in Figure 2.
The atomistic structure of GBs used in this work are shown in Figure 3; the same GBs and interatomic potential were used by Dang et al. [2025]. The atoms in Figure 3 are colored by the Centro Symmetry Parameter (CSP) Kelchner et al. [1998] computed using OVITO Stukowski [2009]. The minimum energy GBs have well-ordered atomic arrangements. Significant deviations from these are clearly present in the higher energy metastable GBs with different microscopic degrees of freedom. To determine the SOECs of the GB region, we performed affine deformation of a bicrystal with 3D periodic boundary conditions. A prescribed set of lattice strains () was applied to an energy minimized bi-crystal under the assumption of general elastic anisotropy. Per-atom energies for the atoms in the GB region were used to calculate the energy density in the GB region as a function of the applied strain. The elastic stiffness coefficients were determined from these energy density-strain data. We assume that the strain in the bi-crystal is homogeneous and equal to the applied strain. The Lamé constants are determined from the SOECs using the Voigt average, see Figure 4.
We perturb the lattice energy in the GB region using an additional term that accounts for the difference in atomic structure and its effect on the evolution of the dislocation core. This additional energy barrier is relevant to the resulting dislocation-GB interaction; however, it cannot be determined analytically but can be estimated from atomic scale simulations. Previous studies have estimated the energy barrier using nudged elastic band (NEB) method [Tsuru et al., 2009] and MD simulations [Sangid et al., 2011, Li et al., 2020, 2022]. In Sangid et al. [2011] and Li et al. [2020], the energy barrier and GBE are correlated using a power law equation; however, the number of data points used are not enough to create a robust model. Thus, in this work, we take a different approach and use the GBE to scale the lattice energy in the bulk grains (see Equation 12) to define the perturbation on the lattice energy in the GB region as
where is a fitting constant, is the GBE informed by MS calculations (See Figure 2), is the unstable stacking fault energy, is the shear modulus of the grain and is the shear modulus of the GB. This term acts as a perturbation of the -surface in the bulk grain to account for the disordered nature of the GB with respect to the bulk crystal so the total energy at the GB region is calculated as . To align with previous work, this term could be modified as where and are fitting parameters [Sangid et al., 2011, Li et al., 2020]. We highlight again that the energy barrier to dislocation glide within a GB due to the atomic disorder is still an open research question, and is expected to have strong dependence on the GB structure. Thus, these equations are representative of this effect, but may need to be redefined for GB structures not considered in this paper. In particular, specialized GB structures such as cube-on-cube orientation or coherent twin boundaries with low values of GBE , are likely examples of cases in which different descriptions of this energy barrier may be necessary.
The different GB widths, GB energies, elastic constants and GB-dislocation reactions as predicted with MD are shown in Table 1. The Lamé constants of the Cu grains surrounding the GB region are and and mJ/m2. For each GB, the minimum energy structure plus two higher metastable structures were selected for this study. We note that we were unable to determine the SOECs for some metastable GBs as they were unstable under some deformation modes. Therefore, the Lamé constants and GB width of the minimum energy GB structure is used also for the metastable cases.
@@
2.4 Material properties and simulation setup
A domain of 256x4x128 is used with a grid size that corresponds to the Burgers vector norm of a screw dislocation in Cu ( Å). Periodic boundary conditions are used in the three directions and the domain size is chosen such that the periodic interactions are minimized, when bigger domains were used the results did not change. An inclusion, representing the GB region, is located in the center of the domain and the position of the initial dislocation dipole is presented in Figure 5. A screw dislocation dipole is placed in the left grain, with the left monopole placed at 63 grid points from the domain boundary with a distance of 50 grid points between the dislocations. The domain is subjected to three different applied shear stresses MPa, resulting in a different reaction for a given GBE value as shown in Table 1. The dislocation mobility is set to with a dimensionless time step of .
3 Results and discussion
The reactions in Table 1 are reproduced using the PFDD model by setting for each type of GB. The values used in the following simulations for the LAGB and HAGB are 0.04 and 0.156, respectively. As the elastic constants and energies are fixed for a given GB, Equation 21 may be redefined as
@@ 3 Results and discussion
The value of is shown in Table 2 for each GB. Based on the reactions at different applied stress, the six GBs can be divided into three categories as shown in Table 2. One can notice the value of is close within each category. If is between 0.1989 and 0.284, the GB-dislocation reactions are (A/P, T, T), if is between 0.656 and 0.743, the GB-dislocation reactions are (A/P, A/P, T), and if is equal or higher than 0.801, the GB-dislocation reactions are (A/P, A/P, A/P). The values roughly correlate with strain-gradient metrics from the strain functional descriptors used in the two-class Gaussian Mixture Model clustering in [Mishra et al., 2024]. More details are provided in the Supplementary Materials, see Supplementary Figures 15-18.
@@
| LAGB | 751 | 0.04 | 0.1989 | A/P | T | T |
| LAGB | 1003 | 0.04 | 0.284 | A/P | T (C1) | T |
| LAGB | 1228 | 0.04 | 0.3605 | A/P | A/P (C2) | T |
| HAGB | 847 | 0.156 | 0.656 | A/P | A/P (C3) | T (C4) |
| HAGB | 938 | 0.156 | 0.743 | A/P | A/P | T |
| HAGB | 998 | 0.156 | 0.801 | A/P | A/P | A/P |
@@
Among the cases shown in Table 1, four different cases are analyzed here in more detail: (C1) LAGB with GBE mJ/m2 under an applied stress of MPa, (C2) LAGB with GBE mJ/m2 under an applied stress of MPa, (C3) HAGB with GBE mJ/m2 under an applied stress of MPa and (C4) HAGB with GBE mJ/m2 under an applied stress of MPa. Note that the other cases have been simulated and the GB-dislocation reactions from MD are reproduced, see Supplementary Figures 1-14. C1, C2 and C3 have different energies and same applied stress while C3 and C4 are the same GB with different applied stresses. for C1, C2 and C3/C4 is 0.284, 0.386 and 0.656, respectively (see Table 2). To the authors knowledge, there is no way to compute analytically the values of , for the work presented here the values are chosen to reproduce the MD reactions.
@@
Figures 6, 7, 8 and 9 show the change of the disregistry and stress evolution at different time steps for cases C1, C2, C3 and C4. The disregistry is computed as [Shen and Wang, 2004]
@@
where is the Burgers vector direction of the initial perfect dislocation. The dislocation transmits in cases C1 and C4, and it is pinned in cases C2 and C3 as the dislocation stays blocked inside (Figure 7) or next (Figure 8) to the GB region. Three main observations need to be highlighted, (i) transmission is easier for LAGB which is in accordance with MD simulations and experimental observations [Nieto-Valeiras et al., 2024, Adams et al., 2019], (ii) the change in GB structure between C1 and C2 is reflected in the increase of GBE, which results in an increased resistance to the dislocation, causing it to be pinned, and (iii) the change in the interaction between cases C3 and C4 from being pinned to transmitted is a direct effect of the applied stress. Also, in C4 the dislocation evolves faster due to the higher applied stress of 750 MPa.
@@
The dislocation disregistry change during interaction with the GB is shown in Figure 10. Moreover, the change in the stacking fault width is clearly shown by plotting its derivative along the slip direction, see Figure 11. These detailed views of the disregistry are shown at the same time steps displayed in Figures 6, 7, 8 and 9. The initial value of the stacking fault width in the grains is 4 or 5 grid points as shown in Figure 11 at t=125, corresponding to 1 nm, which is in agreement with previous studies [Jin et al., 2006]. For the cases where the dislocation transmits across the GB, there is a difference in the stacking fault width during the interaction before and after leaving the GB region. In case C1, the stacking fault width decreases from 4 grid points to 3 grid points before the leading partial enters the GB region, as soon as both the trailing and leading partial enter the GB the stacking fault width decreases to 2 grid points and just before full transmission, the leading partial leaves the GB region generating an increase of the stacking fault width to 6 grid points see Figures 6, 10a and 11. In case C4, the stacking fault width decreases to 2 grid points (Figure 9b), just before full transmission the stacking fault width is 5 grid points and comes back to its initial shape (Figure 9). Something similar happens in the cases that get pinned. In case C2, the dislocation gets pinned with the trailing and leading partials being blocked inside the GB region. Also, the stacking fault width changes slightly to 2 grid points as shown in Figures 10b and 11. In case C3, as the dislocation approaches the GB and gets pinned the stacking fault width decreases to half of its original width from 4 to 2 grid points, see Figures 10c and 11. As the perturbation to the lattice energy is higher in C3 than C2, the partial dislocations constricts to almost a perfect dislocation before the leading partial dislocation slightly enters the GB region. Due to this difference, the interaction in case C2 may be categorized as absorption instead of pinning. Note that absorption is a challenging interaction to model as the interaction can have different scenarios, e.g., dislocations can glide along the GB and re-emit elsewhere, or the dislocation can change the GB structure leading to nucleation of new dislocations at locations along the GB [Jin et al., 2006, Dang et al., 2024, Liang et al., 2019, Wang, 2015]. However, modeling absorption and subsequent dislocation glide along the GB plane is not available in the present PFDD model, but could be considered by adding existing glide and cross slip PFDD extensions [Fey et al., 2022, Wise et al., 2025].
@@
In summary, the variation in the stacking fault width is higher in C3 and C4 as shown in Figures 10 and 11. As is added to the total energy, it acts as an additional attractive force between the two partial dislocations which explains the change in the disregistry. Hence, the difference between the stacking fault width in C1, C2, C3 and C4 is proportional to despite of the higher applied stress in C4. In the Supplementary Material, is plotted against averaged strain gradient metrics showing good correlation with both, net strain gradient and gradient on uniaxial strain metrics computed using MS [Mishra et al., 2024]. Similar changes in the stacking fault width have been reported in MD simulations by Jin et al. [2006] (cf. Figure 2), where the partial dislocation is constricted to an almost perfect dislocation at the GB before transmission or absorption of the partial dislocation, and by Dang et al. [2024] (cf. Figure 2), where the partial dislocation stacking fault width reduces to almost half of its original width when the dislocation is pinned.
The remaining simulations are presented in the Supplementary Material, Supplementary Figures 1-14 show the disregistry and stress evolution, and they are in accordance with the MD reactions as reported in Table 2. Another important observation is that the constriction of the stacking fault width is proportional to the applied stress. This can be seen in the HAGB cases with a GBE of 938 and 998 mJ/m2. When the applied stress is 250 MPa, the partial dislocations get pinned at the GB, and the stacking fault width is about 3 grid points (See Supplementary Figures 9 and 12). When the applied stress is increased to 500 MPa, the partials constrict to almost a perfect dislocation (See Supplementary Figures 10 and 13). At 750 MPa, the LAGB with the highest value of GBE (1228 mJ/m2) transmits as shown in Supplementary Figure 7 while the HAGB with highest GBE (998 mJ/m2) is pinned as shown in Supplementary Figure 14. Even if the GBE of the LAGB is higher, the perturbation to the lattice energy is higher for the HAGB resulting from a combination of the fitting constant, GBE, and the elastic constants. This is clearly evident from the values of in Table 2. This higher perturbation to the lattice energy results in a higher barrier to dislocation transmission in the GB region and consequently higher stress for transmission. Note that the elastic constants of the HAGB and the stress at the GB region is higher than the LAGB, this may affect the reaction. A similar observation was also reported in a recent MD study by Wu et al. [2026].
3.1 Considering multiple GBs
One of the main advantages of mesoscopic simulations is length and time scales can be larger than those considered in MD simulations. In this section, an example of a larger scale simulation is shown where multiple GBs are placed along the domain. The domain size is 729x2x128 grid points with a grid size corresponding to a Burgers vector and periodic boundary conditions in the three directions. Two LAGBs and two HAGBs are placed along the direction, with properties corresponding to the minimum energy state in Table 1. A screw dislocation dipole is placed in the center of the domain, with a distance between the dislocations of 100 grid points and slip direction , see Figure 12. The domain is subjected to a shear stress MPa .
Figure 13 shows the stress field as the dislocations evolve. Each dislocation dissociates into two partials and starts to evolve towards a GB. Due to the different GB properties, the partials interact differently with the GBs, the positive partial dislocations (right) get pinned at the HAGB and the negative partial dislocations (left) get transmitted to the adjacent grain at the early stage of the simulation (See Figure 13a, b and c). The negative partial dislocations continue to evolve to the left (Figure 13d) until it interacts with the next HAGB, constricts and get pinned (Figure 13e and f). This simulation shows the versatility of the new coupled PFDD approach; it allows the model to consider the interaction of dislocations with multiple GBs and can be used to study interaction of more complex GB networks as those present in polycrystals.
@@ 4 Conclusion
A new PFDD formulation that couples phase field and micromechanical solvers is presented. This new formulation allows for the study of heterogeneous materials with different stiffness tensors by solving the Lippmann-Schwinger equation. The total strain obtained from the Lippmann-Schwinger equation is one of the terms in the elastic energy density used within the time-dependent Ginzburg-Landau kinetic equation.
In this work, GB-dislocation interactions are studied by representing GBs as inclusions with a specific width obtained from MS simulations [Mishra et al., 2024]. Two additional parameters are added to describe GBs, (i) an elastic energy which is considered by computing a different GB elastic stiffness tensor and (ii) a lattice energy that represents the effect of the disordered atomic structure within the GB region. Two different groups of GBs are studied including their minimum energy and metastable states: LAGB with GB plane (1,1,6) and misorientation angle of 27∘ and HAGB with GB plane (5,5,2) and misorientation angle of 178∘. The GB-dislocations interactions predicted using PFDD show excellent agreement with the MD results. However, these results depend on the calibration parameter , which varies as a function of the misorientation. We presented an analysis on the stacking fault width, the initial width and reduction of the width during the interaction which shows good agreement with MD simulations [Jin et al., 2006, Dang et al., 2024]. Finally, by adding an additional lattice energy for the GB region proportional to , which correlated with strain gradient metrics derived from atomistic simulations, we can account for the dependency of GB-dislocation interaction on the GB energy or GB structure.
In contrast to the sharp interface method [Zeng et al., 2016, Ma et al., 2022], this new methodology opens the door to model more general GB-dislocation interactions. Future work will be focused on the improvement of the GB description and the inclusion of other interactions such as GB absorption and glide of the dislocation along the GB. Also, other GB-dislocation interactions such as reflection should be considered. Finally, after the model is calibrated using MD data, it is possible to study GB-dislocation interactions in bigger systems such as polycrystals or potentially GBs with varying structure along their length.
Supplementary material
The supplementary material presents the basic scheme algorithm, details of the coupled formulation, the remaining PFDD simulations and plots of strain functional descriptors as a function of